# Enhancing Few-Shot Image Classification through Learnable Multi-Scale Embedding and Attention Mechanisms

### v1

@@ 3.2.3 Distance metric
Once we have obtained the final output from the previous stage, we take the average of the vectors from all the support samples belonging to the same class to obtain the prototypes for each class from the support set as equations 9 .
@@ 4.1 Datasets
The proposed method was evaluated on three widely used datasets commonly employed for few-shot learning tasks: MiniImageNet bib34 , FC100 bib35 , and CUB wah2011caltech .
Table 1: Evaluation on MiniImageNet in 5-way.
Method Backbone 1-shot 5-shot
AdaResNet bib37 ResNet12 56.88 71.94
TADAM bib38 ResNet12 58.50 76.70
MetaOptNet bib39 ResNet12 62.64 78.63
Neg-Margin bib40 ResNet12 63.85 81.57
MixtFSL bib41 ResNet12 63.98 82.04
Meta-Baseline bib42 ResNet12 63.17 79.26
FRN bib43 ResNet12 66.45 82.83
QSFormer bib44 ResNet12 65.24 79.96
Distill bib45 ResNet12 64.82 82.14
Meta-Baseline + DiffKendall bib51 ResNet12 65.56 80.79
LEO bib46 WRN-28-10 61.76 77.59
CC+rot bib47 WRN-28-10 62.93 79.87
FEAT bib48 WRN-28-10 65.10 81.11
Neg-Cosine bib49 ResNet18 62.33 80.94
MixtFSL bib50 ResNet18 60.11 77.76
MergeNet-Concat bib52 ResNet18 65.05 77.76
Our model ResNet18 66.57 84.42
@@
CUB dataset. The CUB (Caltech-UCSD Birds-200-2011) dataset is a collection of bird images often used for fine-grained visual classification tasks. This dataset contains 200 different bird species, with a total of 11,788 images. The images are randomly partitioned into 100 training classes, 50 validation classes, and 50 test classes. This dataset poses a challenging problem as many bird species can appear visually similar, requiring models to learn discriminative features to accurately classify the different species.
Table 2: Evaluation on FC100 in 5-way.
Method Backbone 1-shot 5-shot
P-Net bib30 ResNet12 37.80 53.30
Cosine Classifier bib53 ResNet12 38.47 57.67
TADAM bib54 ResNet12 40.10 56.10
SimpleShot bib55 ResNet10 40.13 53.63
Metaopt Net bib56 ResNet12 41.10 55.50
DC bib57 ResNet12 42.04 57.63
MDM-Net bib58 ResNet12 43.62 57.41
RFS bib60 ResNet12 44.60 60.90
SSFormers bib59 ResNet12 43.72 58.92
LSFSL bib65 ResNet12 43.60 60.12
Barlow Twins + DSA bib66 ViT-B 41.42 55.47
Our model ResNet18 44.78 66.27
Table 3: The comparison of our model’s performance against state-of-the-art methods on the test domain CUB.
Method 1-shot 5-shot
RelatonNet bib61 35.21 51.10
MatchingNet bib62 42.28 57.21
RelationNet+LFT bib63 48.10 65.02
MatchingNet+LFT bib63 43.38 61.44
RelationNet+ATA bib64 48.49 59.42
StyleAdv fu2023styleadv 48.49 68.72
StyleAdv-FT fu2023styleadv - 70.90
Our model 52.95 71.59
@@ 4.2 Experimental setting
For our datasets, we performed the following preprocessing and model configuration steps. First, all input images were resized to a resolution of 84 x 84 pixels. Regarding the model architecture, we employed a pre-trained ResNet-18 as the backbone feature extractor. Specifically, we extracted features from the 5 stages of the ResNet-18 network, obtaining feature maps from each of these stages.
For the optimization of our models, we employed the Adam optimizer. For the MiniImageNet dataset, we used a constant learning rate of $10^{-4}$, while for the FC100 dataset, the learning rate was $2\times 10^{-5}$. This learning rate was selected based on preliminary experiments to ensure stable and efficient convergence of the models during training. All experiments were conducted on an NVIDIA RTX 4090 GPU system.
@@ 4.3 Evaluation metric
baseline Multiscale Learnable Weight Self- attention MiniImageNet FC100
1 shot 5 shot 1 shot 5 shot
X X X 62.83 82.38 39.47 63.44
X X 63.65 83.23 41.97 64.8
X 64.73 83.92 43.7 65.76
66.57 84.42 44.78 66.27
@@ 4.5 Cross domain
To further assess the generalization capabilities of our model, we conducted a cross-domain evaluation using the CUB dataset as the testing domain, while the MiniImageNet dataset was used for training. As shown in Table 3, our model demonstrated strong performance in this challenging cross-domain few-shot learning setting. Specifically, our model achieved an impressive 1-shot accuracy of 52.95 and a 5-shot accuracy of 71.59 on the CUB dataset. This outperforms various state-of-the-art approaches, such as RelationNet+LFT, MatchingNet+LFT, and the strong Baseline model, showcasing the superior ability of our approach to adapt to the distributional shift between the training and testing domains. The consistent improvement over the competing methods across both the 1-shot and 5-shot settings highlights the robustness and transferability of the representations learned by our model. This demonstrates its effectiveness in tackling the challenging cross-domain few-shot learning problem, where the model must generalize its knowledge from the training domain to the novel testing domain. These results further solidify the position of our model as a highly competitive and versatile few-shot learning solution.
@@ 4.6 Ablation study
2) Impact of Weight Initialization: The choice of weight initialization also plays a crucial role in model performance. The best results were observed with initial weights set to 1.1, 1.2, 1.3, and 1.4. This variation underscores the importance of proper weight initialization in achieving optimal model accuracy.
@@ 5 Conclusion
This paper presents an innovative strategy to enhance few-shot classification by integrating a self-attention network and embedding learnable weights at each stage, leading to improved performance and substantial outcomes. Through feature vector extraction and weight transfer across stages, this technique elevates multi-scale feature representation, thereby enhancing overall performance. Moreover, the incorporation of self-attention mechanisms refines features at each stage, resulting in more robust representations and enhanced performance. The introduction of learnable weights at each stage significantly boosts performance and outcomes. Extensive evaluations on the MiniImageNet and FC100 datasets underscore the efficacy of this method compared to current state-of-the-art approaches.

### v2

@@ 3.2.3 Distance metric
Once we have obtained the final output from the previous stage, we take the average of the vectors from all the support samples belonging to the same class to obtain the prototypes for each class from the support set as Equation 8.
@@ 4.1 Datasets
The proposed method was evaluated on three widely used datasets commonly employed for few-shot learning tasks: MiniImageNet bib34 and FC100 bib35 , as well as eight datasets from the CD-FSL benchmark, including TierImageNet bib80 , CUB wah2011caltech , ChestX-ray bib81 , ISIC bib82 , Flower102 bib84 , EuroSAT (Euro) bib83 , CropDisease (CropD) bib85 , and histopathological image dataset bolhasani2020histopathological .
@@ 4.2 Experimental Setting
We performed standard preprocessing and model configuration steps for our datasets. All input images were resized to $84\times 84$ pixels and normalized using a standard normalization technique to improve model convergence. We employed a pre-trained ResNet-18 as the backbone feature extractor and utilized the Adam optimizer for model training. Learning rates and other specific settings are summarized in Table 1.To ensure the reliability of our results, the code was run three times, and the result was reported. All experiments were conducted on an NVIDIA RTX 4090 GPU system.
Table 1: Experimental Settings
Parameter Configuration
Input Image Resolution $84\times 84$ pixels
Model Architecture Pre-trained ResNet-18
Feature Extraction Feature maps from 5 stages
Optimizer Adam
Learning Rate (MinImageNet) $1\times 10^{-4}$
Learning Rate (FC100) $2\times 10^{-5}$
Hardware NVIDIA RTX 4090 GPU
@@ 4.3 Evaluation metric
Table 2: Evaluation on MiniImageNet in 5-way.
Method Year Backbone 1-shot 5-shot
AdaResNet bib37 2018 ResNet12 56.88 71.94
TADAM bib38 2018 ResNet12 58.50 76.70
LEO bib46 2018 WRN-28-10 61.76 77.59
MetaOptNet bib39 2019 ResNet12 62.64 78.63
CC+rot bib47 2019 WRN-28-10 62.93 79.87
Neg-Cosine bib49 2020 ResNet18 62.33 80.94
MixtFSL bib50 2020 ResNet18 60.11 77.76
FEAT bib48 2020 WRN-28-10 65.10 81.11
Neg-Margin bib40 2020 ResNet12 63.85 81.57
Distill bib45 2020 ResNet12 64.82 82.14
FRN bib43 2021 ResNet12 66.45 82.83
MixtFSL bib41 2021 ResNet12 63.98 82.04
Meta-Baseline bib42 2021 ResNet12 63.17 79.26
MergeNet-Concat bib52 2021 ResNet18 65.05 77.76
ViTFSL-Baseline bib67 2022 Transformer 63.51 80.30
QSFormer bib44 2023 ResNet12 65.24 79.96
Meta-Baseline + DiffKendall bib51 2023 ResNet12 65.56 80.79
Auto-MS bib70 2024 HCE-64F 53.33 69.64
Our model 2024 ResNet18 66.57 84.42
Table 3: Evaluation on FC100 in 5-way.
Method Year Backbone 1-shot 5-shot
P-Net bib30 2017 ResNet12 37.80 53.30
TADAM bib54 2018 ResNet12 40.10 56.10
Cosine Classifier bib53 2019 ResNet12 38.47 57.67
SimpleShot bib55 2019 ResNet10 40.13 53.63
Metaopt Net bib56 2019 ResNet12 41.10 55.50
DC bib57 2019 ResNet12 42.04 57.63
RFS bib60 2020 ResNet12 44.60 60.90
MDM-Net bib58 2022 ResNet12 43.62 57.41
SSFormers bib59 2023 ResNet12 43.72 58.92
LSFSL bib65 2023 ResNet12 43.60 60.12
Barlow Twins + DSA bib66 2024 ViT-B 41.42 55.47
Our model 2024 ResNet18 44.78 66.27
Table 4: The comparison of our model’s performance against state-of-the-art methods on CUB and TieredImageNet.
Method TieredImageNet CUB
1-shot 5-shot 1-shot 5-shot
MAML bib20 51.61 65.76 40.51 53.09
ANIL bib72 52.82 66.52 41.12 55.82
BOIL bib76 53.23 69.37 44.20 60.92
Sparse-MAML bib77 53.47 68.83 41.37 60.58
Sparse-ReLU-MAML bib77 53.77 68.12 42.89 57.33
Sparse-MAML+ bib77 53.91 69.92 43.43 62.02
GAP bib78 58.56 72.82 44.74 64.88
Our model 67.40 85.67 52.95 71.59
Table 5: The comparison of our model’s performance against state-of-the-art methods on the test domains of selected datasets in the 5-way 5-shot task.
Method ChessX CropD Euro Flower ISIC
BL++ bib53 25.49 48.22 75.79 63.16 40.73
ANIL bib72 24.41 48.69 63.96 61.27 37.57
ANIL+MLP bib73 25.02 58.10 75.73 66.45 39.19
MTL bib28 24.15 33.27 54.27 58.18 35.56
MTL+MLP bib73 25.19 51.23 65.19 51.34 34.74
PN bib30 24.05 55.59 70.96 63.56 32.95
DN4 bib74 27.34 53.62 75.01 75.01 40.15
CAN bib75 27.46 67.26 78.22 79.81 42.73
Our model 26.33 83.48 78.63 85.13 43.02
Table 6: The comparison of our model’s performance on histopathological image dataset bolhasani2020histopathological in 3-way 1-shot and 3-way 5-shot tasks.
Method 1-shot 5-shot
Baseline 38.42 42.65
Proposed method 39.17 43.94
@@
Baseline Multiscale Learnable Weight Self- attention MiniImageNet FC100 Additional Metrics
1 shot 5 shot 1 shot 5 shot timeInference numParameter
X X X 62.83 82.38 39.47 63.44 0.145s 11M
X X 63.65 83.23 41.97 64.80 0.377s 11M
X 64.73 83.92 43.70 65.76 0.380s 11M
66.57 84.42 44.78 66.27 0.438s 12M
@@ 4.5 Cross domain
To further evaluate the generalization capabilities of our proposed model, we conducted a cross-domain evaluation using the MiniImageNet dataset for training and eight medical datasets as testing domains. These datasets span various medical imaging tasks, providing a comprehensive benchmark for assessing how well models trained on general-purpose datasets can generalize to new and specialized domains.
Table 5 compares our model’s performance with state-of-the-art methods, including BL++, ANIL, CAN, DN4, and MTL+MLP, on challenging 5-way 5-shot tasks. Table 4 further shows performance comparisons on widely used datasets such as CUB and TieredImageNet. Our model demonstrated strong generalization capabilities, outperforming or matching baseline methods in multiple cases.
Additionally, we evaluated our proposed method on a histopathological image dataset bolhasani2020histopathological , which contains three grades and thus represents a 3-way classification problem. Table 6 summarizes the results of this comparison for both 1-shot and 5-shot tasks.
The results indicate that our method successfully adapted to diverse testing domains, showing good performance in tasks such as skin lesion classification, flower identification, and crop disease detection. These findings underscore the effectiveness of our approach in handling domain shifts and limited labeled data.
@@ 4.6 Ablation study
In addition to the improved accuracy, we evaluated the efficiency of the proposed model by comparing the number of parameters and inference time with the baseline. As shown in Table 7, the final model includes one million additional parameters compared to the baseline. Despite this slight increase, the accuracy shows significant improvements across all tasks. Furthermore, the inference time of the final model is approximately three times that of the baseline, which remains a reasonable trade-off considering the substantial accuracy gains. These results highlight the effectiveness of our approach in achieving enhanced performance while maintaining computational efficiency.
@@
2) The choice of weight initialization plays a crucial role in model performance. To investigate this, we started with equal initial weights of 1,1,1,1,1 for all feature vectors. From this starting point, we incrementally increased the weight of each feature vector by 0.1 relative to the previous stage. Among all these configurations, the best results were observed with initial weights of 1,1.1,1.2,1.3,1.4. This process underscores the importance of proper weight initialization, demonstrating that small and systematic adjustments to initial weights can lead to significant improvements in model performance. Furthermore, as shown in the final two rows of Table 8, we also experimented with assigning lower initial weights to deep layers compared to shallow layers. This approach resulted in lower accuracy than configurations where deep layers were assigned higher weights. These findings suggest that deep layers play a more significant and impactful role in classification tasks compared to shallow layers.
@@
4.7 Analysis
In this section, we present a detailed evaluation of the model’s predictions on the MiniImageNet dataset, incorporating both quantitative and qualitative analyses. Visual examples are used to highlight correctly classified and misclassified samples, offering insights into the model’s strengths and limitations in decision-making. Additionally, these visualizations help uncover the underlying factors contributing to correct classifications and prediction errors.
Figure 4 displays correctly classified samples. For example, in the first row, despite similarities between the query image of a vase and support images such as cups, the model successfully classifies the vase with a confidence of 64%. This demonstrates the model’s capacity to distinguish fine-grained details despite visual similarities. Additionally, the green-highlighted prediction scores underscore the model’s robust generalization across challenging query-support pairs.
Conversely, Figure 5 illustrates instances where the model makes incorrect predictions. Upon examination, these misclassifications reveal key challenges in visual reasoning:
  • In the third row, a black-and-white spotted dog in the query image is misclassified as a different breed of dog, likely due to their visual resemblance.
  • Similarly, in other cases, objects with overlapping features or ambiguous contexts appear to confuse the model, leading to reasonable but incorrect predictions.
These observations demonstrate that while the model effectively handles complex cases in many instances, future improvements could further enhance its ability to distinguish visually similar categories and reduce context-driven errors. Addressing these issues is crucial for achieving more accurate classification performance under challenging conditions.
Figure 6 illustrates the confusion matrix for our model on the MiniImageNet dataset, highlighting the class-wise prediction performance and potential misclassification patterns. The results reveal that the model generally distinguishes classes well, but certain visually similar categories present challenges.
For instance, the golden retriever and dalmatian classes show occasional confusion due to shared visual characteristics such as similar body structures and fur patterns. Similarly, the electric guitar class is sometimes misclassified as curiass, likely due to overlapping visual features such as elongated shapes and complex backgrounds. These cases reflect reasonable misclassifications given the inherent visual similarities between query and support images. Despite these challenges, the overall distribution of correct classifications demonstrates the robustness of the model across diverse categories.
Figure 4: Examples of correctly classified samples on the MiniImageNet dataset.
Figure 5: Examples of misclassified samples on the MiniImageNet dataset.
Figure 6: Confusion matrix for the model’s predictions on the MiniImageNet dataset for the 5-way 1-shot task.
@@ 5 Conclusion
This paper presents an innovative strategy to enhance few-shot classification by integrating a self-attention network and embedding learnable weights at each stage, leading to improved performance and significant outcomes. By employing feature vector extraction and weight transfer across stages, our approach elevates multi-scale feature representation, resulting in enhanced overall model performance. The incorporation of self-attention mechanisms effectively refines features at each stage, yielding more robust representations. Extensive evaluations on the MiniImageNet and FC100 datasets demonstrate the efficacy of our method compared to current state-of-the-art approaches. To further validate our model’s generalization capabilities, we conducted experiments across eight cross-domain datasets.Future work will focus on exploring the theoretical rationale behind weight initialization for each stage, which is crucial for optimizing model performance. We also propose a two-phase training approach that eliminates less relevant support images during the initial training phase, allowing the model to concentrate on those most closely related to the query image, ultimately enhancing performance.Furthermore, our method can be adapted and extended to other few-shot learning tasks beyond image classification, such as few-shot object detection and segmentation. To handle larger and more complex datasets, modifications may include refining attention mechanisms and optimizing weight initialization strategies to accommodate increased variability and complexity in the data.